# iTransformer: Inverted Transformers Are Effective for Time Series Forecasting

### 01 Dec 2023 v2

@@ Abstract
The recent boom of linear forecasting models questions the ongoing passion for architectural modifications of Transformer-based forecasters. These forecasters leverage Transformers to model the global dependencies over temporal tokens of time series, with each token formed by multiple variates of the same timestamp. However, Transformers are challenged in forecasting series with larger lookback windows due to performance degradation and computation explosion. Besides, the embedding for each temporal token fuses multiple variates that represent potential delayed events and distinct physical measurements, which may fail in learning variate-centric representations and result in meaningless attention maps. In this work, we reflect on the competent duties of Transformer components and repurpose the Transformer architecture without any modification to the basic components. We propose iTransformer that simply applies the attention and feed-forward network on the inverted dimensions. Specifically, the time points of individual series are embedded into variate tokens which are utilized by the attention mechanism to capture multivariate correlations; meanwhile, the feed-forward network is applied for each variate token to learn nonlinear representations. The iTransformer model achieves state-of-the-art on challenging real-world datasets, which further empowers the Transformer family with promoted performance, generalization ability across different variates, and better utilization of arbitrary lookback windows, making it a nice alternative as the fundamental backbone of time series forecasting.
@@ 1 Introduction
Transformer (Vaswani et al., 2017) has achieved tremendous success in natural language processing (Brown et al., 2020) and computer vision (Dosovitskiy et al., 2021), growing into the foundation model that follows the scaling law (Kaplan et al., 2020). Inspired by the immense success in extensive fields, Transformer with strong capabilities of depicting pairwise dependencies and extracting multi-level representations in sequences is emerging in time series forecasting (Li et al., 2021; Wu et al., 2021; Nie et al., 2023).
@@ 3.1 Structure Overview
$\mathbf{H}^{l+1}$ $=\operatorname{TrmBlock}(\mathbf{H}^{l}),\ l=0,\cdots,L-1,$
@@
where $\mathbf{H}=\{\mathbf{h}_{1},\cdots,\mathbf{h}_{N}\}\in\mathbb{R}^{N\times D}$ contains $N$ embedded tokens of dimension ${D}$ and the superscript denotes the layer index. $\operatorname{Embedding}:\mathbb{R}^{T}\mapsto\mathbb{R}^{D}$ and $\operatorname{Projection}:\mathbb{R}^{D}\mapsto\mathbb{R}^{S}$ are both implemented by multi-layer perceptron (MLP). The obtained variate tokens interact with each other by self-attention and are independently processed by the shared feed-forward network in each $\operatorname{TrmBlock}$. Specifically, as the order of sequence is implicitly stored in the neuron permutation of the feed-forward network, the position embedding in the vanilla Transformer is no longer needed here.
@@ 3.2 Inverted Transformer Components
$\operatorname{LayerNorm}(\mathbf{H})=\left\{\frac{\mathbf{h}_{n}-\operatorname{Mean}(\mathbf{h}_{n})}{\sqrt{\operatorname{Var}(\mathbf{h}_{n})}}\bigg{|}\ n=1,\cdots,N\right\}$ (2)
@@ A.2 Implementation Details
5:$\textbf{for}\ l\ \textbf{in}\ \{1,\cdots,L\}\textbf{:}$$\triangleright$ Run through iTransformer blocks.

### 09 Mar 2024 v3

@@ Abstract
The recent boom of linear forecasting models questions the ongoing passion for architectural modifications of Transformer-based forecasters. These forecasters leverage Transformers to model the global dependencies over temporal tokens of time series, with each token formed by multiple variates of the same timestamp. However, Transformers are challenged in forecasting series with larger lookback windows due to performance degradation and computation explosion. Besides, the embedding for each temporal token fuses multiple variates that represent potential delayed events and distinct physical measurements, which may fail in learning variate-centric representations and result in meaningless attention maps. In this work, we reflect on the competent duties of Transformer components and repurpose the Transformer architecture without any modification to the basic components. We propose iTransformer that simply applies the attention and feed-forward network on the inverted dimensions. Specifically, the time points of individual series are embedded into variate tokens which are utilized by the attention mechanism to capture multivariate correlations; meanwhile, the feed-forward network is applied for each variate token to learn nonlinear representations. The iTransformer model achieves state-of-the-art on challenging real-world datasets, which further empowers the Transformer family with promoted performance, generalization ability across different variates, and better utilization of arbitrary lookback windows, making it a nice alternative as the fundamental backbone of time series forecasting. Code is available at this repository: https://github.com/thuml/iTransformer.
@@ 1 Introduction
Transformer (Vaswani et al., 2017) has achieved tremendous success in natural language processing (Brown et al., 2020) and computer vision (Dosovitskiy et al., 2021), growing into the foundation model that follows the scaling law (Kaplan et al., 2020). Inspired by the immense success in extensive fields, Transformer with strong capabilities of depicting pairwise dependencies and extracting multi-level representations in sequences is emerging in time series forecasting (Wu et al., 2021; Nie et al., 2023).
@@ 3.1 Structure Overview
$\mathbf{H}^{l+1}$ $=\operatorname{TrmBlock}(\mathbf{H}^{l}),\ l=0,\dots,L-1,$
@@
where $\mathbf{H}=\{\mathbf{h}_{1},\dots,\mathbf{h}_{N}\}\in\mathbb{R}^{N\times D}$ contains $N$ embedded tokens of dimension ${D}$ and the superscript denotes the layer index. $\operatorname{Embedding}:\mathbb{R}^{T}\mapsto\mathbb{R}^{D}$ and $\operatorname{Projection}:\mathbb{R}^{D}\mapsto\mathbb{R}^{S}$ are both implemented by multi-layer perceptron (MLP). The obtained variate tokens interact with each other by self-attention and are independently processed by the shared feed-forward network in each $\operatorname{TrmBlock}$. Specifically, as the order of sequence is implicitly stored in the neuron permutation of the feed-forward network, the position embedding in the vanilla Transformer is no longer needed here.
@@ 3.2 Inverted Transformer Components
$\operatorname{LayerNorm}(\mathbf{H})=\left\{\frac{\mathbf{h}_{n}-\operatorname{Mean}(\mathbf{h}_{n})}{\sqrt{\operatorname{Var}(\mathbf{h}_{n})}}\bigg{|}\ n=1,\dots,N\right\}$ (2)
@@ 5 Conclusion and Future Work
6 Ethics Statement
Our work only focuses on the time series forecasting problem, so there is no potential ethical risk.
7 Reproducibility Statement
In the main text, we have strictly formalized the model architecture with equations. All the implementation details are included in the Appendix, including dataset descriptions, metrics, model, and experiment configurations. The code will be made public once the paper is accepted.
Acknowledgments
This work was supported by the National Key Research and Development Plan (2021YFB1715200), the National Natural Science Foundation of China (U2342217 and 62022050), the BNRist Innovation Fund (BNR2024RC01010), the CCF-Ant Green Computing Fund, and the National Engineering Research Center for Big Data Software.
@@ A.2 Implementation Details
5:$\textbf{for}\ l\ \textbf{in}\ \{1,\dots,L\}\textbf{:}$$\triangleright$ Run through iTransformer blocks.